OpenStudy (anonymous):

differentiate using logarithmic differentiation y=(2(x^2+1))/((cos2x)^(1/2))

OpenStudy (zarkon):

take the log of both sides...simplify...differentiate

OpenStudy (anonymous):

whoops 2 on the outside because its an exponent

OpenStudy (anonymous):

use the log property: lna -lnb=ln a/b

OpenStudy (anonymous):

so would it just be the natural log of the top - the natural log of the bottom?

OpenStudy (zarkon):

you can simplify more than that

OpenStudy (anonymous):

do you mind explaining? i was recently put into calc 2 skipping calc 1 completely, a little confused =/

OpenStudy (zarkon):

\[\ln(a^r)=r\ln(a)\]

OpenStudy (zarkon):

\[\ln(y)=\ln(2)+\ln(x^2+1)-\frac{1}{2}\ln(\cos2x)\]

OpenStudy (anonymous):

ohhh alright makes a lot more sense now! thank you!

OpenStudy (anonymous):

now differentiate........1/y dy/dx=0+1/(x^2+1) *2x -1/2 1/cos2x * -2sin2x

OpenStudy (anonymous):

back sub your y and multiply it over simplify and your good

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

is the co2x and the -2sin2x both in the denom? or is the -2sin2x inthe numerator.

OpenStudy (zarkon):

\[\frac{1}{\cos(2x)}{(-2\sin(2x))}\]

OpenStudy (anonymous):

alright gotcha, thank you